OpenStudy (anonymous):

just need pointed in the right direction here as far as the first step in the equation... seven fourths decreased by a number is 2. find the number do i subtract seven fourths from both sides or what? im so lost...

OpenStudy (anonymous):

Write out the word problem as an equation. What do you get?

OpenStudy (mertsj):

Start with the 7/4. Write it down. Now, how do you decrease something?

OpenStudy (anonymous):

7/4-n=2

OpenStudy (mertsj):

Perfect!!

OpenStudy (anonymous):

do i add n to both sides and then solve or do i subbtract 7/4 from both sides?

OpenStudy (anonymous):

Both will eventually get you the same result.

OpenStudy (mertsj):

Depends on whether you prefer the coefficient of the variable to be positive or not. On this one, I would probably subtract the 7/4 and then change the answer to its opposite. In other words, multiple both sides by -1

OpenStudy (mertsj):

*multiply not multiple

OpenStudy (anonymous):

ok so 7/4-n=2 -n=2-7/4 -n=2-13/4 -n=1/4 -n*-1=1/4*-1 n=-1/4 ???

OpenStudy (anonymous):

1 and 3/4 oops

OpenStudy (mertsj):

Be careful. Remember 2 = 8/4

OpenStudy (anonymous):

FYI, if you use the equation button below you can use frac{}{} to write out fractions.

OpenStudy (anonymous):

oh cool, thanks im new to this...

OpenStudy (mertsj):

Yes. n= -1/4

OpenStudy (anonymous):

yay! maybe im getting the hang of this lol

OpenStudy (anonymous):

Place your value for n into the original equation and solve it. \[\frac{7}{4}-n=2\] \[\frac{7}{4}--\frac{1}{4}=2\] \[\frac{7}{4}+\frac{1}{4}=2\] \[\frac{8}{4}=2\]

OpenStudy (anonymous):

thanks, again for all yalls help!